OpenStudy (anonymous):

A species population dropped from 9,000 in the year 2,000 to 4,500 in the year 2011. Find the percent decrease in it's population . HELP PLEASE

ganeshie8 (ganeshie8):

so whats the difference in population ?

OpenStudy (aa123):

hi

OpenStudy (aa123):

hey @ganeshie8 can u help me with my question after this

OpenStudy (anonymous):

4500 ?

ganeshie8 (ganeshie8):

yes! compare it with the starting population of 9000

OpenStudy (anonymous):

What do you mean compare ?

ganeshie8 (ganeshie8):

take the percent : percent decrease = \[\large \dfrac{\text{difference }}{\text{original value}}\times 100\]

OpenStudy (anonymous):

oh okay

ganeshie8 (ganeshie8):

percent decrease = \[\large \dfrac{\text{4500}}{\text{9000}}\times 100\]

OpenStudy (anonymous):

50